# Rethinking Probabilistic Circuit Parameter Learning

### v1

@@ 3.2 Extension to the Mini-Batch Case
When the dataset is large, full-batch EM becomes inefficient and impractical as it requires scanning the entire dataset before making any parameter updates. In such cases, we instead wish to update the parameters after processing only a small subset of data points, which is commonly referred to as a mini-batch. Given a mini-batch of samples $\mathcal{D}$, Peharz et al. [2020] proposes to update the parameters according to a small step size $\alpha\in(0,1)$:
$\theta^{\prime}_{n,c}=(1-\alpha)\cdot\theta_{n,c}+\alpha\cdot\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)/Z,$ (5)
where we borrow notation from Equation 4. However, it remains unclear whether this update rule is grounded in a principled EM objective. In the following, we show that from the full-batch EM derivation in the previous section, we can derive a mini-batch update rule that closely resembles the above, but with a crucial difference.
@@
To account for this, we can put a weighting factor $\gamma>1$ on the KL divergence (i.e., $\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$ becomes $\gamma\cdot\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$). [^5] Note that this is equivalent to $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})-(\gamma-1)\cdot\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$ according to Proposition 1. Plugging in Lemma 1 and dropping terms independent to $\boldsymbol{\phi}^{\prime}$, the adjusted objective (i.e., $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})$ with the additional weighting $\gamma$) can be expressed as
@@
where $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}$ collects all entries $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)$ into a single vector, with each $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)$ representing the aggregated gradient w.r.t. $\phi_{n,c}$. With the constraints that ensure each PC node defines a normalized distribution, the solution is
$\theta^{\prime}_{n,c}=\left(\mathtt{TD}_{\boldsymbol{\phi}}(n)\cdot\theta_{n,c}+\alpha\cdot F_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)\right)/Z,$ (6)
where $\alpha:=1/(\gamma-1)$ is the learning rate, $\mathtt{TD}_{\boldsymbol{\phi}}(n)$ is the TD-prob of node $n$ (Def. 2), and $Z$ is a normalizing constant. The derivation is deferred to Section B.2. In practice, compared to the full-batch EM update (Eq. (4)), the only additional computation required is $\mathtt{TD}_{\boldsymbol{\phi}}(n)$, which can be efficiently computed once in a data-independent top-down pass over the PC using Definition 2.
@@
where we define $\log\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})$ as the (unnormalized) log-likelihood of node $n$. A key observation is that the second term $\hat{\mathtt{F}}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)$ is normalized w.r.t. all children of sum node $n$: $\sum_{c\in\mathsf{ch}(n)}\hat{\mathtt{F}}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)=1$ (see Appx. B.3 for the proof). Intuitively, we now break down $\mathtt{F}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)$ into the importance of node $n$ to the overall output (i.e., $\mathtt{F}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n)$) and the relative contribution of child $c$ to $n$ (i.e., $\hat{\mathtt{F}}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)$). With this decomposition, we can simplify Equation 6 as
$\theta^{\prime}_{n,c}=\big{(}\theta_{n,c}+\alpha\cdot\mathtt{rel}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n)\cdot\hat{\mathtt{F}}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)\big{)}/Z,$
where $\mathtt{rel}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n):=\mathtt{F}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n)/\mathtt{TD}_{\boldsymbol{\phi}}(n)$ can be viewed as the relative importance of $n$ to the PC output given input $\boldsymbol{x}$. The term $\alpha\cdot\mathtt{rel}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n)$ then acts as an adaptive learning rate for updating the child parameters of node $n$, scaling the update magnitude according to how influential $n$ is for the input. In comparison, the mini-batch algorithm in Equation 5 uses a fixed learning rate for all parameters.
This difference is reflected in the example shown in Figure 1. Given the PC on the left, which represents a mixture of two Gaussians (middle). Suppose we draw one sample $x\!=\!-1.5$. This sample does not reflect the full distribution and only activates the left mode. Our algorithm accounts for this by assigning a small effective learning rate to node $n_{2}$ ($\mathtt{rel}_{\boldsymbol{\phi}}^{x}(n_{2})\!\approx\!0.0004$) as it is “not responsible” for explaining this particular input and focuses the update on $n_{1}$ ($\mathtt{rel}_{\boldsymbol{\phi}}^{x}(n_{1})\!\approx\!1.9996$). In contrast, Equation 5 applies a uniform learning rate across all parameters, leading it to also update $n_{2}$ unnecessarily to fit the current sample.
Figure 1: The proposed algorithm implicitly applies an adaptive learning rate to each node. For the PC shown on the left, given a sample $x\!=\!-1.5$, the algorithm uses a large learning rate to update $n_{1}$ while keeping $n_{2}$ almost unchanged.
@@ 4 Connections with Gradient-Based Optimizers
Solving for $\boldsymbol{\phi}^{\prime}$ gives $\boldsymbol{\phi}^{\prime}=\boldsymbol{\phi}+\alpha\cdot\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}$, a standard gradient ascent step with $\alpha=1/(2\gamma)$.
We empirically compare our mini-batch EM algorithm against existing EM- and gradient-based optimizers on training a PC to fit the ImageNet32 dataset [Deng et al., 2009], where the proposed algorithm converges faster and also achieves higher likelihoods after convergence (Appx. D).
Acknowledgements.

### v2

@@ 3.2 Extension to the Mini-Batch Case
When the dataset is large, full-batch EM becomes inefficient and impractical as it requires scanning the entire dataset before making any parameter updates. In such cases, we instead wish to update the parameters after processing only a small subset of data points, which is commonly referred to as a mini-batch. Given a mini-batch of samples $\mathcal{D}$, Peharz et al., (2020) propose to update the parameters following:
$\theta^{\prime}_{n,c}=(1-\alpha)\cdot\theta_{n,c}+\alpha\cdot\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)/Z,\;(\alpha\in(0,1))$ (6)
where we borrow notation from Equation 5. However, it remains unclear whether this update rule is grounded in a principled EM objective. In the following, we show that from the full-batch EM derivation, we can derive a mini-batch update rule that closely resembles the above, but with a crucial difference.
@@
To account for this, we can put a weighting factor $\gamma>1$ on the KL divergence (i.e., $\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$ becomes $\gamma\cdot\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$). [^7] Note that this is equivalent to $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})-(\gamma-1)\cdot\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$ according to Proposition 1. Plugging in Lemma 1 and dropping terms independent to $\boldsymbol{\phi}^{\prime}$, the adjusted objective (i.e., $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})$ with the additional weighting $\gamma$) becomes
@@
where $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}$ collects all entries $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)$ into a single vector, with each $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)$ representing the aggregated gradient w.r.t. $\phi_{n,c}$. With the constraints that ensure each PC node defines a normalized distribution (i.e., for each sum node $n$, $\sum_{c\in\mathsf{ch}(n)}\theta^{\prime}_{n,c}\!=\!1$), the solution is
$\theta^{\prime}_{n,c}=\left(\mathtt{TD}_{\boldsymbol{\phi}}(n)\cdot\theta_{n,c}+\eta\cdot F_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)\right)/Z,$ (7)
where $\eta:=1/(\gamma-1)$ is the learning rate, $\mathtt{TD}_{\boldsymbol{\phi}}(n)$ is the TD-prob of node $n$ (Def. 2), and $Z$ is a normalizing constant. The derivation is deferred to Section B.2. In practice, compared to the full-batch EM update (Eq. (5)), the only additional computation required is $\mathtt{TD}_{\boldsymbol{\phi}}(n)$, which can be efficiently implemented using any autodiff library to compute the gradient of $Z(\boldsymbol{\phi})$ w.r.t. the log-parameters $\phi_{n,c}$ (see proof in Appx. B.4).
@@
where we define $\log\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})$ as the (unnormalized) log-likelihood of node $n$. A key observation is that the second term $\hat{\mathtt{F}}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)$ is normalized w.r.t. all children of sum node $n$: $\sum_{c\in\mathsf{ch}(n)}\hat{\mathtt{F}}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)=1$ (see Appx. B.3 for the proof). Intuitively, we now break down $\mathtt{F}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)$ into the importance of node $n$ to the overall output (i.e., $\mathtt{F}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n)$) and the relative contribution of child $c$ to $n$ (i.e., $\hat{\mathtt{F}}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)$). This simplifies Equation 7 as
$\theta^{\prime}_{n,c}=\big(\theta_{n,c}+\eta\cdot\mathtt{rel}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n)\cdot\hat{\mathtt{F}}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n,c)\big)/Z,$
where $\mathtt{rel}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n):=\mathtt{F}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n)/\mathtt{TD}_{\boldsymbol{\phi}}(n)$ can be viewed as the relative importance of $n$ to the PC output given $\boldsymbol{x}$. The term $\eta\cdot\mathtt{rel}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(n)$ then acts as an adaptive learning rate for updating the child parameters of $n$, scaling the update magnitude according to how influential $n$ is for $\boldsymbol{x}$. In comparison, the mini-batch algorithm in Equation 6 uses a fixed learning rate for all parameters.
This difference is reflected in the example shown in Figure 1. Given the PC on the left, which represents a mixture of two Gaussians (middle). Suppose we draw one sample x=1.5x\!=\!-1.5. This sample does not reflect the full distribution and only activates the left mode. Our algorithm accounts for this by assigning a small effective learning rate to node $n_{2}$ ($\mathtt{rel}_{\boldsymbol{\phi}}^{x}(n_{2})\!\approx\!0.0004$) as it is “not responsible” for explaining this particular input and focuses the update on $n_{1}$ ($\mathtt{rel}_{\boldsymbol{\phi}}^{x}(n_{1})\!\approx\!1.9996$). In contrast, Equation 6 applies a uniform learning rate across all parameters, leading it to also update $n_{2}$ unnecessarily to fit the current sample.
Figure 1: The proposed algorithm implicitly applies an adaptive learning rate to each node. For the PC shown on the left, given a sample x=1.5x\!=\!-1.5, the algorithm uses a large learning rate to update $n_{1}$ while keeping $n_{2}$ almost unchanged.
@@ 4 Connections with Gradient-Based Optimizers
Solving for $\boldsymbol{\phi}^{\prime}$ gives $\boldsymbol{\phi}^{\prime}=\boldsymbol{\phi}+\eta\cdot\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}$, a standard gradient ascent step with $\eta=1/(2\gamma)$.
The Regularization Terms.   While this analogy highlights a shared structure between EM and gradient-based optimizers, it also reveals a fundamental discrepancy. The KL divergence in the EM formulation is a natural measure of distance between distributions, while the L2 penalty in gradient-based updates only constrains the movement in parameter space. This distinction is important because proximity in parameter space does not necessarily translate to proximity in distribution space. For instance, adding all parameters $\boldsymbol{\phi}$ (note that they represent log-probabilities) by a constant leaves the distribution unchanged, yet the L2 penalty would still register this as a large deviation. Therefore, compared to gradient-based methods, the EM formulation better respects the geometry of distributions.
Normalization Constraints.   Another notable difference between gradient-based methods and the EM algorithms (both the full-batch and the mini-batch ones) is the treatment of normalization constraints. In EM, parameters are updated in a way that preserves local normalization, i.e., the edge parameters of each sum node $n$ are guaranteed to sum to one after every update. On the other hand, standard gradient-based optimization does not enforce this constraint, and thus intermediate parameter values may fall outside the normalized parameter space.
One might naturally wonder whether this discrepancy leads to better or worse training dynamics, since optimization would be carried out in an enlarged parameter space. Perhaps surprisingly, we show that it has no effect on the optimization process, as the parameters $\boldsymbol{\phi}$ can always be mapped to a locally normalized counterpart $\boldsymbol{\phi}$’ that preserves both the represented distribution and the gradients with respect to the parameters.
Specifically, given a PC with parameters $\boldsymbol{\phi}$ (corresponds to $\boldsymbol{\theta}:=\exp(\boldsymbol{\phi})$), there exists a normalization algorithm that outputs $\boldsymbol{\phi}^{\prime}$ and ensures (i) all parameters are locally normalized (i.e., $\forall n,\sum_{c\in\mathsf{ch}(n)}\theta_{n,c}=1$), (ii) the represented distributions are unchanged (i.e., $\forall\boldsymbol{x},\tilde{{p}}_{\boldsymbol{\phi}}(\boldsymbol{x})\propto\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}(\boldsymbol{x})$), and (iii) the gradients with respect to the parameters are preserved (i.e., $\forall\boldsymbol{x},\partial\tilde{{p}}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}=\partial\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}(\boldsymbol{x})/\partial\boldsymbol{\phi}^{\prime}$.
Intuitively, the above three conditions guarantee that applying this normalization algorithm neither alters the model’s probabilistic semantics nor interferes with the optimization dynamics. In other words, the model continues to represent the same distribution, and the gradients used for subsequent updates remain the same. Therefore, it can be applied after each gradient update without changing the learning dynamics.
In the following, we present the normalization algorithm and show how it can be implemented efficiently. [^8] The existence of this normalization algorithm has been previously shown by Martens and Medabalimi, (2014), although they did not provide an explicit algorithm. A detailed analysis and formal proofs of the three properties are provided in Appendix C.
Figure 2: Log-Likelihood over epochs on four diverse datasets. For the ImageNet (YCC-R) and ImageNet (YCC) datasets, an HCLT with hidden size 512 is used; for the WikiText dataset, an HMM with hidden size 256 is used; for the BioBank dataset, a PDHCLT with hidden size 1024 is used. Anemone achieves significantly faster convergence as well as final $L$ across all four cases.
Table 1: Negative $L$s on the UK BioBank Chromosome 6 dataset. Anemone consistently and significantly outperforms all baseline across every tested model architecture. The results also show a clear performance hierarchy among the baseline methods. The best result in each column is marked in bold.
Optimizer BioBank Chr6
HCLT 512 HCLT 1024 PDHCLT 512 PDHCLT 1024
Full-batch EM 55.3 53.8 46.5 45.1
Adam 102.7 100.4 112.4 110.4
Mini-batch EM 55.7 55.5 49.5 47.2
Anemone (Ours) 54.5 52.1 45.3 42.2
Table 2: Convergence speed (epochs) for PDHCLT 1024 on BioBank Chr6 Dataset. The table reports epochs to reach specific $L$ thresholds, with $\Delta$ representing the difference from the best $L$ of -45.1 before anemone (Table 1). Lower is better. Bold marks the best result per column; $\infty$ indicates failure to reach the threshold in time.
Method $L$ $\geq$ -48 $L$ $\geq$ -46 $L$ $\geq$ -45.1
($\Delta\approx 2.9$) ($\Delta\approx 0.9$) ($\Delta=0$)
Full EM 450 645 1000
Adam $\infty$ $\infty$ $\infty$
Mini EM 715 $\infty$ $\infty$
Anemone 50 80 130
The Algorithm.   The normalization procedure consists of two simple passes over the PC. First, we perform a feedforward evaluation of the PC to compute the partition function $Z_{n}(\boldsymbol{\phi})$ at every node $n$:
[…]
Next, for every edge $(n,c)$ where $n$ is a sum node and $c\in\mathsf{ch}(n)$, we update the parameter as
$\phi^{\prime}_{n,c}=\log\left(\frac{\theta_{n,c}\cdot Z_{\boldsymbol{\phi}}(c)}{Z_{\boldsymbol{\phi}}(n)}\right).$ (8)
5 Experiments
In this section, we empirically evaluate anemone against existing EM- and gradient-based optimizers across a range of PC models and datasets. Section 5.1 introduces the models, datasets, and baseline optimizers. In Section 5.2, we ask whether anemone can achieve higher log-likelihoods at convergence compared to existing approaches, and whether anemone converges faster. Finally, Section 5.3 investigates the key design factors in setting hyperparameters through a series of ablation studies.
5.1 Experimental Setup and Baselines
We conduct our empirical evaluation across three distinct domains (i.e., DNA sequence, image, and text) using various PC architectures.
DNA Sequence Modeling.
We evaluate density estimation performance on a high-dimensional genomics dataset from the UK Biobank (Sudlow et al.,, 2015). For this task, we train Hidden Chow-Liu Trees (HCLTs) (Liu and Van den Broeck,, 2021) and their variant that combines the PD structure (Poon and Domingos,, 2011) termed Partitioned Data HCLTs (PDHCLTs), which is defined in Appendix E. A detailed description of the dataset and all model configurations is provided in Section D.1.
Image Modeling.   We adopt the ImageNet32 dataset (Deng et al.,, 2009) with two color transformations, i.e., a lossy YCC transformation and its lossless variant YCC-R (Malvar and Sullivan,, 2003). Each $\times32$ image is partitioned into four $\times16$ image patches, which results in a total of $16\times 16\times 3=768$ variables. We employ HCLTs and PDHCLTs for both datasets. Details of the datasets and the models are included in Section D.1.
Language Modeling.   We use the WikiText-103 (Merity et al.,, 2017) dataset, which is widely used for language modeling. The dataset is preprocessed by the GPT-2 tokenizer (Radford et al., 2019a, ). We evaluate the Hidden Markov Model (HMM) PC architecture and its recently proposed variant Monarch HMM (Zhang et al.,, 2025) on it.
EM-Based Baselines.   We adopt two EM baselines, which are the standard full-batch EM and the mini-batch EM proposed by Peharz et al., (2020) (cf. Eq. (6)). For the full-batch EM, there are no hyperparameters to tune. In contrast, for the mini-batch EM, we tune the batch size in the range $\{512,16384\}$ and the step size $\alpha$ in Eq. (6) over $\{0.05,0.1,0.4\}$.
Table 3: Negative $L$s on the ImageNet32 dataset’s validation subset. Anemone consistently outperforms other baselines when training HCLT models. We observe a general performance ranking where Adam optimizer is the weakest, followed by Mini-batch EM and Full-batch EM. OOM denotes out-of-memory errors due to computing resource limitations. The best result in each column is marked in bold.
Optimizer ImageNet32 YCC-R ImageNet32 YCC
PDHCLT 256 HCLT 512 HCLT 1024 HCLT 512 HCLT 1024
Full-batch EM 2529 2480 2469 2164 2163
Adam 2553 2518 OOM 2187 OOM
Mini-batch EM 2529 2506 2470 2179 2232
Anemone (Ours) 2530 2477 2469 2158 2159
Table 4: Negative $L$s on the WikiText-103 dataset. Anemone achieves the best $L$s on three of the four tested model configurations and is highly competitive on the fourth. The best results are marked in bold.
Optimizer WikiText
HMM 256 HMM 512 HMM 1024 Monarch HMM 1024
Full-batch EM 722.6 702.2 682.8 738.1
Adam 735.7 OOM OOM OOM
Mini-batch EM 725.2 703.2 682.1 734.6
Anemone (Ours) 722.2 701.7 682.3 734.0
Gradient-Based Baselines.   We adopt the Adam optimizer (Adam et al.,, 2014), which is used by many recent works. We tune the batch size and the learning rate in the ranges $\{512,1024\}$ and $\{1\times 10^{-2},3\times 10^{-3},5\times 10^{-3},1\times 10^{-3}\}$, respectively.
Our Method.   For ease of definition, we express the step size of our method as $\alpha=\eta/(\eta-1)$, where $\eta$ is given in Equation 7. The hyperparameter search is detailed in Section D.2. Similar to the mini-batch EM algorithm, we use the proposed momentum update with $\beta=0.9$. We propose to apply a momentum update to the flows $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)$. Specifically, we initialize the momentum flows $\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)=\mathbf{0}$, then before every EM step, we update $\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)$ following:
$\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)\leftarrow\beta\cdot\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)+(1-\beta)\cdot\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c).$
Finally, we replace the $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)$ in Equation 6 with $\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)/(1-\beta^{T+1})$, where $T$ is the number of updates performed. We compare the performance with and without the momentum update for both anemone and mini-batch EM in Section 5.3.
5.2 Overall Performance and Convergence
We begin by examining the training dynamics of different optimizers. Figure 2 shows the training curves (i.e., validation LL vs. number of epochs) across all four datasets, each paired with an appropriate PC architecture (see the caption for the details).
We start by focusing on the three baseline approaches, i.e., full-batch EM, Adam, and mini-batch EM. A consistent trend is that full-batch EM consistently reaches better (or comparable) performance at convergence. This is a strong indicator that the existing mini-batch optimizers (Adam and mini-batch EM) fail to compensate for the reduced information available in a mini-batch compared to the full dataset.
In contrast, despite being a mini-batch algorithm, anemone achieves significantly better performance at convergence, even compared to full-batch EM. We evaluate the final log-likelihoods of each optimizer across a wider range of PC architectures. The final log-likelihoods, evaluated either at convergence or after a fixed maximum number of epochs, are shown in Tables 1, 3 and 4 for the DNA, image, and text modeling tasks, respectively. On the BioBank Chr6 dataset, anemone achieves consistent and significant performance gains over all baselines. For the ImageNet32 and WikiText datasets, it again obtains better results in the majority of cases.
Additionally, as indicated by Figure 2, anemone converges faster than all baselines, including the mini-batch ones that are designed for faster convergence. We further quantify the number of epochs used to reach a certain LL. Specifically, as shown in Table 2 anemone requires $\sim$8x fewer epochs to reach the same validation LL. Further experiments are deferred to Section D.4.
5.3 Ablation Study
To disentangle the performance gains of anemone from the general effects of momentum (cf. Table 4), we conduct ablation study shown in Figure 3. While the results confirm that momentum improves the final log-likelihood for anemone (left panel), they also show that it provides little benefit when applied to vanilla mini-batch EM (right panel), suggesting that the improvements are not from momentum alone, but from the synergy between momentum and anemone, that is not observed in standard mini-batch approaches.
Figure 3: Ablation study on the effect of momentum when combined with anemone and vanilla mini-batch EM, respectively. Left: For anemone optimizer (HMM with hidden size of 256 on WikiText), incorporating momentum improves the final log-likelihood despite slightly slower initial convergence, while still being significantly faster than full-batch EM. Right: In contrast, for mini-batch EM (PDHCLT with hidden size of 512 on BioBank Chr6), momentum provides little benefit.
6 Related Works and Conclusion
Modeling Advancements in PCs.   The development of PCs has been marked by significant progress in enhancing their expressiveness and utility as generative models. Since the initial establishment of PCs, research has focused on developing more expressive and scalable PC structures. Specifically, a line of research has sought to design PC structures that are expressive yet parameter-efficient (Rahman et al.,, 2014; Adel et al.,, 2015; Liu and Van den Broeck,, 2021), while another set of approaches pursues iterative structure learning strategies that progressively expand the model’s capacity (Liang et al.,, 2017; Dang et al.,, 2022; Di Mauro et al.,, 2021; Liu et al., 2023b, ). Both directions have contributed to significant performance gains on widely used text and image datasets.
Parameter Learning of PCs.   Beyond structure, a central challenge in learning PCs lies in the optimization of their parameters. This problem has been studied from both the algorithmic and the systems perspective. On the algorithmic side, two families of approaches dominate: EM-style updates and gradient-based methods. The EM algorithm was first applied to PCs by Poon and Domingos, (2011), and later extended to mini-batch training in Peharz et al., (2020). Gradient-based optimizers such as Adam (Kingma,, 2014) have also become a common choice in practice due to their simplicity and scalability.
On the systems side, considerable effort has gone into developing efficient implementations that can handle the large computational and memory demands of PCs. Optimized einsum backends (Peharz et al.,, 2020), specialized libraries such as SPFlow (Molina et al.,, 2019), and more recently PyJuice (Liu et al.,, 2024) provide high-performance primitives that enable scaling to PCs with billions of parameters.
Conclusion.   This work addresses the underperformance of mini-batch optimizers for PCs. We identify that existing methods ineffectively regularize distribution changes, causing them to “overfit” to the current mini-batch. We propose anemone, a novel mini-batch EM algorithm that applies an implicit adaptive learning rate to each parameter, scaled by its contribution to the batch likelihood. Across extensive experiments, anemone consistently outperforms existing optimizers in both convergence speed and final performance.
Acknowledgements